### v3

# Policy Learning with Confidence

Abstract.
This paper introduces a rule for policy selection in the presence of estimation uncertainty, explicitly accounting for estimation risk. The rule belongs to the class of risk-aware rules on the efficient decision frontier, characterized as policies offering maximal estimated welfare for a given level of estimation risk. Among this class, the proposed rule is chosen to provide a reporting guarantee, ensuring that the welfare delivered exceeds a threshold with a pre-specified confidence level. We apply this approach to the allocation of a limited budget among social programs using estimates of their marginal value of public funds and associated standard errors.
This is a revised version of the arXiv working paper 2502.10653 originally posted in February 2025. We thank Isaiah Andrews, Jiafeng (Kevin) Chen, Nathan Hendren, Toru Kitagawa, Matt Masten, Adam McCloskey, Whitney Newey, Vira Semenova, Jesse Shapiro, Aleksey Tetenov and conference and seminar participants at the Munich Econometrics Workshop, Duke University, CUNY, University of Geneva, Tilburg University, the Encounters in Econometric Theory workshop, CUHK Shenzhen, IAAE, Goethe University Frankfurt, the 2025 ESWC, and the 2026 ASSA meeting for helpful comments and discussion. Tian Xie provided excellent research assistance. Liyang Sun acknowledges generous support from the Economic and Social Research Council (new investigator grant UKRI607).
Department of Economics, MIT
Department of Economics, Columbia University.
§Department of Economics, Duke University.
Department of Economics, UCL and CEMFI. Email: liyang.sun@ucl.ac.uk
Keywords: budget allocation, risk-aware policy learning, statistical decision theory
JEL classification codes: C14, C44, C52.
1. Introduction
Consider a decision maker (DM) faced with choosing from a menu of policies to maximize expected welfare. The DM could be a planner deciding how to allocate treatments to heterogeneous individuals, a firm choosing different potential innovations in which to invest resources, an auctioneer choosing an optimal auction design or reserve price, or a policy-maker choosing expenditure shares to allocate to government programs to maximize a measure of public welfare. If the DM knew the welfare that would be obtained from implementing each policy, the choice would be easy; she would then simply choose the policy that would yield the highest welfare. However, the DM lacks such precise knowledge, but instead has data from one or possibly multiple studies that can be used to consistently estimate the welfare that would be achieved by each policy. The estimates are measured with varying degrees of precision, as reflected by their standard errors. How should the DM use the available information for policy selection?
An intuitively appealing choice is to select the policy with the highest estimated welfare, the so-called plug-in rule, which simply replaces the population objective with its sample analog in determining policy choice. In the literature on policy learning, this is referred to as empirical welfare maximization (EWM). [^1] See for example Kitagawa and Tetenov (2018) and Athey and Wager (2021). Andrews et al. (2024) alternatively refers to this as the “natural rule” or “picking the winners”. The same logical task applies more broadly to a variety of economic contexts, for example when researchers use structural models, they obtain estimates of the performance of different policies that can then be used to inform policy choice. Manski (2021) calls this “as-if” optimization, described as, “specification of a model, point estimation of its parameters, and use of the point estimate to make a decision that would be optimal if the estimate were accurate.” Here we refer to such rules interchangeably as plug-in rules or empirical welfare maximization (EWM) rules.
Unfortunately, the available estimates for the welfare of different choices are generally not identical to their population values. Estimation error could result in the EWM rule delivering suboptimal policy choice. It may be observed, for example, that among a menu of options policy A has the highest estimated welfare with policy B coming in a close second, but that the standard error of policy B is much smaller than that of policy A, reflecting that the welfare of policy B is estimated much more precisely. Could it be that policy B is actually better than policy A, and that the higher estimated welfare of policy A is simply down to a lack of precision in the estimates? Should the DM consider choosing policy B since it is estimated more precisely to hedge against estimation error?
This paper proposes a rule for selecting policies with the goal of maximizing expected welfare in the presence of estimation uncertainty, explicitly accounting for estimation risk in the decision rule. As inputs to her decision the DM has consistent estimates for the welfare of each available policy, and their corresponding sample variance matrix, with sample estimates approximately normally distributed around their population means. [^2] If the estimates are from independent samples the sample variance matrix will simply be a diagonal matrix with each policy’s sample variance along the diagonal. We define and analyze a class of risk-aware policy rules that provide a principled manner for explicitly balancing the estimated welfare of each policy against its estimation risk, and which we show have favorable regret properties. Risk-aware rules select a policy from the (Pareto) efficient decision frontier that balances performance and precision as measured by trading off higher sample estimates against smaller sample variance, where the exact tradeoff is governed by the DM’s choice of policy rule from the class of risk-aware rules. Our newly proposed rule is determined from within this class by balancing this tradeoff using the tangent line to the frontier whose slope is determined by the critical value of a one-sided upper confidence band for welfare following the approach for intersection bound inference developed in Chernozhukov et al. (2013). Consequently, this rule delivers a reporting guarantee, ensuring with high confidence that the actual welfare delivered exceeds a lower threshold, while also having the favorable regret properties of all risk-aware rules. The proposed rule thus ensures Policy Learning with Confidence, and is subsequently referred to as the PoLeCe rule.
To illustrate, consider a setting in which a budget-conscious DM needs to allocate funds across several social programs that affect the same group of individuals, either in response to a slight budget increase or for incremental cost-cutting. Suppose the DM wishes to use the Marginal Value of Public Funds (MVPF), spearheaded by Hendren and Sprung-Keyser (2020), for this purpose. The MVPF of each social program is the ratio of its marginal benefit to its net marginal cost to the government, inclusive of the impact of any behavioral responses on the government budget. For instance, an MVPF of 1.5 indicates that each \$1 of net government spending generates \$1.50 in benefits for beneficiaries. For a utilitarian DM, the MVPF provides a metric to compare the “bang for the buck” of different policies that target the same beneficiaries. A DM interested in allocating additional marginal expenditure will prioritize programs with the highest MVPF to maximize welfare. [^3] Here we abstract from distributional incidence and assume the DM values dollars equally across the beneficiaries of each policy. If welfare weights vary across beneficiaries, as in the original framework of Hendren and Sprung-Keyser (2020), our analysis can accommodate this by adjusting each MVPF in the value function using the DM’s designated weights. Example welfare weights have been derived in Hendren (2020). Other metrics for measuring programs’ marginal benefits and costs have also been discussed in García and Heckman (2022).
In practice the DM must make decisions based on estimated MVPFs. The Policy Impacts Library, Hendren and Sprung-Keyser (2021), provides easily accessible MVPF estimates online while also noting significant estimation uncertainty in some cases. Using the EWM rule to allocate funds would direct the entire budget to the program with the highest MVPF estimate. However, the DM may be reluctant to direct the entire budget to a single social program if this MVPF estimate has a large standard error relative to other programs whose MVPF estimates are nearly as high, and which have smaller standard errors. How should the DM balance the tradeoff between the size of the estimated MVPF and the precision with which it is estimated?
[FIGURE 1 AROUND HERE]
This tradeoff is illustrated in Figure 1 using the MVPF estimates from the Policy Impacts Library, which comprises 14 US programs after we focus on estimates based on randomized control trials (RCTs) with finite standard errors. To mimic a DM who applies constant welfare weights within certain age groups but not necessarily across them due to distributional concerns, we use the Policy Impacts Library classification of programs based on intended beneficiaries’ age. We then consider budget allocations separately for two distinct age groups. Six of these programs target beneficiaries age 25 and under, covering early childhood education programs, college financial aid, and job training. [^4] As explained in the Policy Impacts Library, alternative specifications have led to larger MVPF estimates for early childhood education depending on how lifetime earnings are forecasted and whether one accounts for the transfer value of preschool subsidies to parents. Eight of these programs target beneficiaries age 25 and above, covering federal social assistance, health insurance, and housing vouchers. In both cases, the MVPF of the programs selected by EWM are not estimated as precisely as the welfare delivered by diversifying the budget across several programs whose MVPFs are estimated with significantly higher precision. If a DM values both welfare and precision, how should these goals be balanced?
The PoLeCe rule is determined by the intersection of the green line in each panel of Figure 1 with the decision frontier for RW-rules. The rule maximizes estimated welfare offset by a data-dependent penalization factor such that the resulting objective value automatically provides the lower bound of a one-sided confidence interval for the welfare achieved by the selected rule. Thus, the rule delivers a reporting guarantee, ensuring with high confidence that the actual welfare exceeds a lower threshold; no adjustment for post-selection inference is required. Importantly, policies that allocate fractional shares to different treatments or programs can be allowed, producing a richer frontier than that available from singleton allocations such as the individual programs indicated by hollow squares in Figure 1.
In the above example, all that was required to map out the efficient decision frontier, and indeed all that is needed to determine the optimal allocation following our proposal, are point estimates for the welfare of each available policy and their joint variance-covariance. [^5] Our proposed PoLeCe rule uses the sample correlation of the estimates to calibrate the precision penalty to achieve the aforementioned reporting guarantee. The general class of risk-aware decision rules described in Section 2.1 allows for different choices for the penalization factor, which corresponds to the green slope highlighted in Figure 1. Other risk-aware rules may thus only require standard errors rather than estimates of the entire correlation structure, depending on their choice of penalization factor. This can be obtained in applications that feature a variety of different models and sampling processes. An important special case to which our framework applies is the analysis of optimal treatment assignment using data on individual-specific treatments and allocations, such as from an RCT. This is demonstrated in online Appendix H with three additional applications spanning the study of treatments that encourage immunization, enable informal savings technologies, and encourage sobriety among low-wage workers. In a real-world setting of treatment assignment, while the nonprofit GiveWell uses EWM based on RCTs, they have recently highlighted the importance of accounting for estimation uncertainty in decision-making for the sake of “transparency” (Salisbury, 2024). Our proposal offers a practical solution to this concern, while also applying to settings in which individual-specific treatment responses may not be directly observable, and when estimates may be obtained from different sources.
Prescriptions for decision making with sample data include conditional Bayes rules, maximin rules, and minimax regret rules. These are summarized in Manski (2021), which advocates the use of statistical decision theory from the frequentist perspective of Wald (1950). Minimax regret rules in particular have recently received renewed attention, starting with the pioneering work of Manski (2004). Follow-up work includes a decision theoretic framework introduced by Dehejia (2005), finite-sample bounds considered by Stoye (2009), and an asymptotic framework introduced by Hirano and Porter (2009, 2020). While the use of statistical decision theory for selecting policy rules summarized in Manski (2021, 2023) is conceptually appealing, computational challenges remain, particularly in settings involving samples from multiple data sources as in the preceding example. Recent papers on policy learning provide approaches for treatment rule estimates with favorable asymptotic regret properties. Kitagawa and Tetenov (2018) prove the optimality of EWM rules in the sense that as the sample size increases expected regret converges to zero at the minimax rate. Mbakop and Tabord-Meehan (2021) proposes a penalized welfare maximization rule that penalizes the complexity of each policy, and establishes an oracle property for model selection. Athey and Wager (2021) and Demirer et al. (2019) propose doubly-robust estimation of average welfare, which leads to an optimal rule even with quasi-experimental data. The rule proposed here provides a practical approach to providing a decision rule with favorable regret properties while also yielding a confidence interval for welfare. As such, the PoLeCe rule fits within the class of P-certified decision rules studied by Andrews and Chen (2025), a point we come back to in Section 3.2 where we discuss the comparison to other recently proposed rules and inference approaches.
The rest of the paper proceeds as follows. Section 2.1 sets out the framework and defines the class of risk-aware (RW) decision rules. Section 2.2 establishes high probability regret bounds for RW-rules in the spirit of regret properties analyzed for EWM rules by e.g. Kitagawa and Tetenov (2018) and Athey and Wager (2021). Section 3.1 sets forth the PoLeCe rule among the class of RW decision rules and formally provides its dual regret and coverage guarantees. Section 3.2 discusses the general class of P-certified decisions and other rules and inference approaches proposed in the recent literature. Section 4 returns to the context of a planner deciding on the allocation of additional public funds across different social programs and presents the full application of our approach to this empirical setting. Section 5 concludes. Proofs of propositions and details of the Gaussian bootstrap used to compute the PoLeCe penalization factor are provided in the appendices. Additional results including alternative motivations for risk-aware decisions, details for implementation and construction of the efficient decision frontier, three empirical applications to treatment choice using data from RCTs, and computational experiments calibrated to those applications are included in the online supplementary appendices.
2. Risk-Aware Policy Choices
2.1. Motivation and Risk-Aware Policy Choice
We begin with the following key problem as motivation. A public or private agency is considering programs $\{1,...,J\}$ and must decide on investment shares
$\pi=(\pi_{1},...,\pi_{J}),$
subject to
$\pi\in\Pi:=\Bigl\{\,(\pi_{1},\dots,\pi_{J}):\ \sum_{j=1}^{J}\pi_{j}=1,\ 0\leq a_{j}\leq\pi_{j}\leq b_{j}\leq 1\Bigr\},$
which is the unit simplex intersected with a rectangular set of constraints. These additional constraints may reflect diversity or other requirements. The welfare of allocation $\pi$ is
$V(\pi)=\pi^{\prime}R,$
where $R$ is a vector of the rate of return measures, for instance the ratios of marginal benefit to net cost of funds used in public finance. The agency is given estimates $\widehat{R}$ that are approximately Gaussian [^6] In the empirical application we consider, $\Omega$ is block-diagonal, and $n$ is the notional sample size used to study the behavior of rules as more information is acquired.
$\widehat{R}\;\overset{a}{\sim}\;N(R,\Omega/n),$
and the estimated welfare of the allocations are also approximately Gaussian by the continuous mapping theorem:
$\widehat{V}(\pi)=\pi^{\prime}\widehat{R}\;\overset{a}{\sim}\;\pi^{\prime}R+s(\pi)Z_{\pi},\quad Z_{\pi}\sim N(0,1),\quad s^{2}(\pi)={\pi^{\prime}\Omega\,\pi}/{n},$ (1)
where $\{Z_{\pi}\}_{\pi\in\Pi}$ is a Gaussian vector with standard normal marginals. Here $\overset{a}{\sim}$ means “approximately distributed as” formalized in condition (G) below.
Hence, for each fund allocation $\pi$, the agency is given $\widehat{V}(\pi)$ along with its associated estimation risk $s(\pi)$. What should the agency do? Our proposal is to use the risk-aware rule that maximizes empirical welfare offset by the estimation risk times a critical value:
$\widehat{\pi}_{\text{RW}}\;\in\;\operatorname*{argmax}\Bigl\{\widehat{V}(\pi)-k\,\widehat{s}(\pi):\pi\in\Pi\Bigr\},$ (2)
where $k$ is a critical value and $\widehat{s}(\pi)$ is a consistent estimator of the estimation risk $s(\pi)$. By varying $k$, we trace out the efficient decision frontier of treatment policies, as illustrated by Figure 1 in the introduction. Each point on the frontier corresponds to a particular $k>0$. Appendix G in the online supplement provides an efficient algorithm to compute the frontier for any application.
Our leading proposal to choose $k$ is to meet certain reporting guarantees with high confidence, thereby conducting “policy learning with confidence” (PoLeCe). The resulting choice generally differs from the empirical welfare maximizer (EWM):
$\widehat{\pi}_{\text{EWM}}\;\in\;\operatorname*{argmax}\{\widehat{V}(\pi):\pi\in\Pi\}.$ (3)
In the budget allocation problem, the EWM rule would simply allocate all funds to the program with the highest estimated return. This is generally unappealing on intuitive grounds. In contrast, the risk-aware approach (2) would spread out funds over an “efficient” portfolio of programs, similar to Markowitz portfolio allocation in finance. However, the motivation and formulation of our approach are distinct from the Markowitz model.
2.2. Regret Properties and Guarantees of Risk-Aware Policies
[…]
$\widehat{\pi}_{\text{RW}}(\widehat{k})\;\in\;\arg\max_{\pi\in\Pi}\bigl\{\widehat{V}(\pi)\;-\;\widehat{k}\,\widehat{s}(\pi)\bigr\},$
where $\widehat{k}\geq 0$ may depend on both the decision maker’s preferences and the data. Note that $\widehat{k}=0$ corresponds to the EWM rule.
Define
$\widehat{Z}_{\pi}\;:=\;\frac{\widehat{V}(\pi)-V(\pi)}{\widehat{s}(\pi)}$
to be the normalized estimation error process. We use the following Gaussian approximation condition, denoted by (G). Suppose the policy class $\Pi$ can be well-approximated by a $p$-dimensional discretization. Let $\mathcal{A}$ represent the collection of rectangular sets in $\mathbb{R}^{p}$.
  • (G)
    There exists a sequence of non-negative constants $r_{n}$ with $r_{n}\searrow 0$ such that,
    $\Bigl\lvert\Pr\!\bigl((\widehat{Z}_{\pi})_{\pi\in\Pi}\in A\bigr)\;-\;\Pr\!\bigl((Z_{\pi})_{\pi\in\Pi}\in A\bigr)\Bigr\rvert\;\leq\;r_{n},\quad\text{ for all }A\in\mathcal{A}.$ (G)
These conditions are known to hold under mild assumptions. [^7] See Chernozhukov et al. (2022, 2023a) for sharp forms of Gaussian approximations, and Chernozhukov et al. (2023b) for a review. Additional references, e.g. Belloni et al. (2018a); Quintas-Martinez (2022) discuss Gaussian approximations for many or a continuum of target parameters, including those learned via debiased machine learning. Furthermore,  (G) implicitly requires consistency of risk estimates: $\max_{\pi\in\Pi}\lvert\widehat{s}(\pi)/s(\pi)-1\rvert\to_{P}0$.
Quantiles of the Maximal Estimation Error. For a subset $K\subseteq\Pi$, let
$q_{1-\beta,K}\;:=\;(1-\beta)\text{-Quantile}\Bigl(\sup_{\pi\in K}Z_{\pi}\Bigr)$
denote the $(1-\beta)$-quantile of the estimation error over $K$ under the Gaussian approximation. Under (G), with probability at least $1-\beta-r_{n}$, $\max_{\pi\in K}\widehat{Z}_{\pi}\leq q_{1-\beta,K}$. Let $V_{\max}=\sup_{\pi\in\Pi}V(\pi)$ be the maximal true welfare and $\Pi_{0}=\{\pi\in\Pi:V(\pi)=V_{\max}\}$ be the set of policies with maximal true welfare, with cardinality $p_{0}:=|\Pi_{0}|$. Define
$\bar{\sigma}_{\Pi}\;:=\;\sqrt{n}\,\max_{\pi\in\Pi}\widehat{s}(\pi),\quad\underline{\sigma}_{\Pi_{0}}\;:=\;\sqrt{n}\,\min_{\pi\in\Pi_{0}}\widehat{s}(\pi).$
These represent, respectively, an upper bound on the estimation risk of all policies in $\Pi$ and a lower bound on the estimation risk of the best policies $\Pi_{0}$.
Proposition 1 (Regret Bounds for Risk-Aware Decisions).
[…]
In particular, if $0\leq\widehat{k}\lesssim q_{1-\beta,\Pi}$, then with the same probability,
$V_{\max}-V(\widehat{\pi}_{\text{RW}}(\widehat{k}))\;\;\lesssim\;\;\bigl(\bar{\sigma}_{\Pi}/\sqrt{n}\bigr)\,q_{1-\beta,\Pi}.$
Moreover, if $q_{1-\beta,\Pi}\leq\widehat{k}\lesssim q_{1-\beta,\Pi}$, then with the same probability,
$V_{\max}-V(\widehat{\pi}_{\text{RW}}(\widehat{k}))\;\;\lesssim\;\;\bigl(\underline{\sigma}_{\Pi_{0}}/\sqrt{n}\bigr)\,q_{1-\beta,\Pi}.$
As $\beta$ decreases, the probability that the regret bound (4) of Proposition 1 holds increases, since the failure probability is at most $2\beta+2r_{n}$. However, this comes at the cost of a looser bound because the quantile $q_{1-\beta,\Pi}$ increases as $\beta$ gets smaller. This reflects a basic trade-off: a smaller $\beta$ provides a more conservative bound that holds with higher probability, while a larger $\beta$ results in a tighter bound that is less certain to be valid.
Proposition 1 provides dimension-free bounds. To obtain simpler dimension-based bounds, observe that by standard concentration inequalities, [^8] This follows from Lemma A.8 in Belloni et al. (2018b). Talagrand-style upper bounds can also be used, but for this these simpler concentration bounds suffice.
$q_{1-\beta,K}\;\;\leq\;\;\mathrm{E}\Bigl(\sup_{\pi\in K}Z_{\pi}\Bigr)\;+\;\sqrt{2\log(1/\beta)}\;\;\leq\;\;\sqrt{2\log|K|}\;+\;\sqrt{2\log(1/\beta)}\;=\;u_{1-\beta,|K|}.$
We now see that the broad class of RW-rules has regret bounded by
$\bar{\sigma}_{\Pi}\,\sqrt{\frac{\log p}{n}},$
as $p\to\infty$ (and $n\to\infty$). This class includes the EWM rule ($k=0$) and matches known optimal minimax rates when available (see Athey and Wager (2021)).
A key observation of Proposition 1 is that choosing $\widehat{k}>0$ sufficiently large can yield a regret bound of
$\underline{\sigma}_{\Pi_{0}}\,\sqrt{\frac{\log p}{n}},$
which has the same rate but may have a much smaller multiplicative constant
$\underline{\sigma}_{\Pi_{0}}<\bar{\sigma}_{\Pi}$
if there is significant heterogeneity in estimation risk across policies. The leading constant in the regret bound is governed by the least risky among the best policies, and can therefore be significantly smaller than that of the EWM rule. The empirical risk minimization literature has also used penalties on estimation risk to improve error bounds, albeit in different contexts; see, for example, Maurer and Pontil (2009); Swaminathan and Joachims (2015); Foster and Syrgkanis (2023).
These ideas naturally lead to the next section, where we propose a main policy rule that carefully constructs a data-driven $\widehat{k}\approx q_{1-\alpha,\Pi}$ to achieve small regret and provide additional reporting guarantees.
3. The PoLeCe Rule
3.1. Reporting Guarantees
In practice, to obtain reporting guarantees for RW decisions, one needs to approximate $q_{1-\alpha,\Pi}$. We do so via the bootstrap, where we set:
$\widehat{q}_{1-\alpha,\Pi}\;:=\;(1-\alpha)\text{-Quantile}\Bigl(\max_{\pi\in\Pi}\widehat{Z}_{\pi}^{*}\Bigr),$
[…]
  • (B)
    With probability at least $1-\delta_{n}$ (where $\delta_{n}\searrow 0$),
    $\Bigl\lvert{\Pr}^{*}\!\bigl((\widehat{Z}_{\pi}^{*})_{\pi\in\Pi}\in A\bigr)\;-\;\Pr\!\bigl((Z_{\pi})_{\pi\in\Pi}\in A\bigr)\Bigr\rvert\;\leq\;r_{n},\quad\text{for all }A\in\mathcal{A}.$ (B)
Like (G), condition (B) is satisfied under mild assumptions even when $p$ is much larger than $n$, and has been verified for a variety of estimation methods, including debiased machine learning. [^9] See Chernozhukov et al. (2022, 2023a) and Chernozhukov et al. (2023b) for more details, along with other references on “approximate means” settings such as debiased ML (Belloni et al. (2018a); Quintas-Martinez (2022)).
A key consequence is that, letting $r_{n}^{\prime}:=2r_{n}+\delta_{n}$, we have
$\sup_{\pi\in\Pi}\widehat{Z}_{\pi}\;\;\leq\;\;\widehat{q}_{1-\alpha,\Pi}\quad\text{with probability at least }1-\alpha-r_{n}^{\prime},$ (5)
(see Lemma 1 for a proof). Inequality (5) then directly implies a uniform lower confidence bound (LCB) on the performance of all policies:
Proposition 2 (LCB Guarantees for RW Decisions).
With probability at least $1-\alpha-r_{n}^{\prime}$, the true value of any policy $\pi\in\Pi$ is bounded below as follows:
$V(\pi)\;\;\geq\;\;LV_{1-\alpha}(\pi)\;:=\;\widehat{V}(\pi)\;-\;\widehat{q}_{1-\alpha,\Pi}\,\widehat{s}(\pi)\ \text{ for all }{\pi\in\Pi}.$
We now introduce policy learning with confidence (PoLeCe), a risk-aware rule that provides a key reporting guarantee by maximizing the LCB on policy welfare:
$\widehat{\pi}_{\mathrm{PoLeCe}}\;\;\in\;\;\operatorname*{argmax}_{\pi\in\Pi}LV_{1-\alpha}(\pi).$ (6)
Notably, PoLeCe is an RW decision that uses the bootstrap quantile
$\widehat{k}=\widehat{q}_{1-\alpha,\Pi}$
as the penalty for estimation risk. Furthermore, the maximized LCB
$LV_{1-\alpha,\Pi}\;:=\;\max_{\pi\in\Pi}LV_{1-\alpha}(\pi)\;=\;\max_{\pi\in\Pi}\Bigl\{\widehat{V}(\pi)\;-\;\widehat{q}_{1-\alpha,\Pi}\,\widehat{s}(\pi)\Bigr\}$
provides a natural performance guarantee since $V\left(\widehat{\pi}_{PoLeCe}\right)\geq LV_{1-\alpha,\Pi}$ with probability at least $1-\alpha-r_{n}^{\prime}$ by Proposition 2.
Remark 1 (PoLeCe as Minimizer of the UCB on Regret).
We can also view $\widehat{\pi}_{\mathrm{PoLeCe}}$ as minimizing an upper confidence bound on regret. Indeed, for any $\pi\in\Pi$,
$\mathrm{Regret}(\pi)\;=\;V_{\max}-V(\pi)\;\;\leq\;\;V_{\max}\;-\;\bigl(\widehat{V}(\pi)-\widehat{k}\,\widehat{s}(\pi)\bigr)$
with probability at least $1-\alpha-r_{n}^{\prime}$. Hence, by its construction, $\widehat{\pi}_{\mathrm{PoLeCe}}$ minimizes this UCB on regret:
$\min_{\pi\in\Pi}\Bigl\{V_{\max}\;-\;\bigl(\widehat{V}(\pi)-\widehat{k}\,\widehat{s}(\pi)\bigr)\Bigr\}.$
The following result establishes formal properties of the PoLeCe rule.
Proposition 3 (Reporting Guarantees and Regret Bounds for PoLeCe).
With probability at least $1-\alpha-r_{n}^{\prime}$,
$V_{\max}\;\;\geq\;\;V(\widehat{\pi}_{\mathrm{PoLeCe}})\;\;\geq\;\;LV_{1-\alpha,\Pi}.$
Moreover, with probability at least $1-\alpha-\beta-r_{n}-r_{n}^{\prime}$, the regret of $\widehat{\pi}_{\mathrm{PoLeCe}}$ satisfies
$V_{\max}-V(\widehat{\pi}_{\mathrm{PoLeCe}})\;\;\leq\;\;V_{\max}-LV_{1-\alpha,\Pi}\;\;\leq\;\;\bigl(\underline{\sigma}_{\Pi_{0}}/\sqrt{n}\bigr)\Bigl\{q_{1-\beta,\Pi_{0}}\;+\;q_{1-\alpha+r_{n},\Pi}\Bigr\},$
where $\underline{\sigma}_{\Pi_{0}}=\sqrt{n}\,\min_{\pi\in\Pi_{0}}\widehat{s}(\pi)$ is determined by the least risky policy among those with maximal true policy welfare.
The first result in Proposition 3 shows that $LV_{1-\alpha,\Pi}$ provides a valid high-confidence lower bound on the true welfare of the PoLeCe rule. The second result sharpens the general regret bounds presented in Proposition 1. As noted following Proposition 1, the parameter $\beta$ governs the trade-off between failure probability and bound tightness, while the constant $\alpha$, chosen by the researcher, determines the confidence level of the lower bound $LV_{1-\alpha,\Pi}$. The refinement in Proposition 3 depends on the relationship between $\alpha$ and $\beta$. When $\alpha=\beta$, Propositions 1 and 3 yield qualitatively identical guarantees. When $\alpha<\beta$, the bounds in both propositions remain qualitatively identical, but the bound in Proposition 3 holds with higher probability. When $\alpha>\beta$, the bound in Proposition 3 is tighter, but it may hold with lower probability. For a large $\alpha$, the bound may fail to hold; in such cases, the general bound from Proposition 1 remains valid and informative, as it applies to all risk-aware decision rules.
3.2. Discussion and Related Literature
The PoLeCe rule maximizes $LV_{1-\alpha}(\pi)$, the welfare estimate of each policy $\pi$, penalized by the critical value $\widehat{q}_{1-\alpha,\Pi}$ times its standard error $\widehat{s}(\pi)$. The function $LV_{1-\alpha}(\cdot)$ provides a $1-\alpha$ lower confidence band for $V(\cdot)$ following the construction of Chernozhukov et al. (2013). Thus, the PoLeCe rule can be obtained by minimizing worst-case loss over a $1-\alpha$ lower confidence band for welfare. As such, it falls within the class of rules Manski (2021) calls “as-if optimization with set estimates”. Section 3.2 of Manski (2021) suggested that one might consider decision making using confidence sets, but left this to future research.
The contemporaneously developed working paper Andrews and Chen (2025) introduces the term P-certified decisions for the class of rules that ensure an upper bound on loss – equivalently a lower bound on welfare – with probability at least $1-\alpha$. Andrews and Chen (2025) argue that such decisions are useful for risk-averse decision makers and show that “as-if” decisions using confidence sets form an essentially complete class of P-certified decisions. A similar finding is made in Kiyani et al. (2025), which studies rules that maximize worst-case loss over conformal prediction sets. Within this class, those that comprise upper contour sets of conditional quantiles of utility are found to be optimal. Use of these sets for as-if optimization with set estimates differs from the use of upper confidence bands that corresponds to the PoLeCe rule. Ben-Michael et al. (2025) also propose a P-certified decision rule for criminal release recommendations. Their rule again differs from ours, as it is based on a two-sided confidence set for expected utility. [^10] The setting studied by Ben-Michael et al. (2025) has other distinguishing features, for example using data based on deterministic functions of pre-trial risk assessment scores, and the need to take on the challenge of partially identified policy welfare. While the PoLeCe rule falls within the class of P-certified decisions studied in Andrews and Chen (2025), its P-certificate is based on confidence bands of the type studied in Chernozhukov et al. (2013) and it thus differs from the other policy rules studied within this class. A convenient feature of the PoLeCe rule is that it is delivered directly by maximizing $LV_{1-\alpha}(\pi)$ in a single stage without requiring explicitly computing a maximin rule over a set estimate. Furthermore, it provides a rule that lies on the efficient decision frontier with the balance between estimated performance and sample variation pinned down by the critical value corresponding to the planner’s desired confidence level.
P-certified decisions are also connected to the recent literature on inference on winners, e.g. Benjamini et al. (2019), Andrews et al. (2024), Zrnic and Fithian (2024, 2025). These approaches provide confidence sets for the welfare of the selected choice that has the highest empirical welfare in-sample. Inference approaches for optimal assignment and/or welfare in the population include those in the online supplement of Kitagawa and Tetenov (2018), Rai (2019), Armstrong and Shen (2023), and Ponomarev and Semenova (2024). Proposition 3 shows that the PoLeCe rule provides a coverage guarantee for both the welfare of the selected policy and the optimal rule.
Recent alternatives that account for sampling uncertainty in policy selection include Sun (2025) and Moon (2025). Sun (2025) considers policy selection when the decision maker faces a budget constraint and there is uncertainty regarding the cost of the policy. Moon (2025) proposes an empirical Bayes approach to deal with statistical uncertainty in the welfare of different policies. The approach in this paper differs from each of these.
4. Empirical Examples of Investment Allocations for Public Programs
This section continues the discussion from the introduction, and illustrates how to solve the problem of investment allocations for public programs described in Section 2.1. We obtain MVPF estimates $\widehat{R}$ from the Policy Impacts Library, Hendren and Sprung-Keyser (2021). There are 172 programs in total, and we focus on those in the United States whose MVPF is estimated by an RCT. We assume the reported upper and lower bounds correspond to 95% confidence intervals and infer the standard errors for each $\widehat{R}$ from these bounds, yielding $14$ programs. [^11] We exclude programs for which the upper and lower bounds are infinite. We also assume the estimation errors across programs are independent, so their covariance matrix is diagonal, with the squared standard errors on the diagonal. Note that here we only account for estimation error, holding the specification in the original papers that produced these MVPF estimates fixed. Alternative specifications can lead to different MVPF estimates for the same program. For example, accounting for transfers to parents could lead to larger MVPF estimates for early childhood programs.
In the following illustrations, we implement PoLeCe by reformulating the problem as a root-finding task that involves solving second-order cone programs, which has superior computational efficiency relative to grid search as detailed in Appendix B. We set $\alpha=0.05$ throughout.
Table 1 reports the allocation selected by PoLeCe and EWM among programs targeting individuals older than 25. As expected, the EWM rule is a corner solution that places all investment in the highest MVPF program according to the empirical estimates, Holistic Wrap-around Services Can Improve Employment Rates. The PoLeCe rule, in contrast to EWM, places weight on Medicaid for single adults (OHIE, Single Adults) and job training for adults (JTPA) as their MVPF estimates are among the highest, and are highly precise.
[TABLE 1 AROUND HERE]
Table 2 reports the allocation selected by PoLeCe and EWM among programs targeting individuals aged 25 and under. Compared to allocations for programs for those over age 25 in Table 1, the PoLeCe solutions here allocate a large share towards programs whose MVPF estimates are less precise. This is because $\widehat{q}_{0.95,\Pi}$ is smaller, resulting in less aversion to estimation uncertainty.
[TABLE 2 AROUND HERE]
5. Conclusion
In this paper we have focused on the problem faced by a DM who wishes to select from a menu of policies to maximize welfare using imperfect sample estimates. In order to balance the estimated performance of each policy with its associated statistical uncertainty, we proposed and analyzed the properties of a class of risk-aware policies that make the precision/performance tradeoff explicit. Such policies achieve favorable regret properties. We proposed a specific rule from the class of risk-aware policies, namely the PoLeCe rule, which uses a data-dependent construction for balancing the inherent tradeoff between estimated performance and sample uncertainty, such that a lower confidence bound on the welfare obtained by the chosen policy is automatically provided.
A large body of work synthesized in Manski (2013, 2019, 2024) has advocated for greater acknowledgement and incorporation of uncertainty in planning problems. This paper contributes by proposing a principled way to acknowledge and incorporate statistical uncertainty into decision making. Statistical uncertainty is however only one of the many different types of uncertainty that may be present. [^12] For example, Manski (2019) discusses transitory uncertainty, permanent uncertainty, and conceptual uncertainty. The statistical uncertainty considered here is one type of permanent uncertainty. We have focused on settings in which the welfare of each policy is point identified, such that consistent estimates of the welfare and sample variance of the policies are available. Application of the concepts developed here to settings that feature so-called “deep uncertainty”, or ambiguity, would seem an important direction for further development given the large number of settings in which the mean performance of various policies is only credibly partially identified. [^13] The literature on treatment choice when mean treatment performance is partially identified goes back to Manski (2000). Recent research with references to the broader literature includes Russell (2020), Ishihara and Kitagawa (2021), Yata (2021), Christensen et al. (2023), Kido (2023), Montiel Olea et al. (2025), and Ben-Michael et al. (2025). This would require balancing statistical uncertainty with a collection of interval estimates for each policy’s performance.
While the nuance required to formally develop such an approach is beyond the scope of the present paper, a rough prescription could be made based on an extension of the reporting guarantee established here. To see how, consider the definition of the PoLeCe rule in (6). The rule selects the policy that maximizes a $1-\alpha$ lower confidence band for the maximal achievable welfare. Construction of such a confidence band does not require that the optimal level of welfare or the optimal rule be point identified, and could be implemented using techniques developed in e.g. Chernozhukov et al. (2013) under partial identification. This is of course not the only way to balance statistical uncertainty and ambiguity. A more thorough study of the performance of such an approach, and the implicit tradeoff between statistical uncertainty and ambiguity due to partial identification would seem a useful direction for future research.
Appendix A Proofs
In what follows let
$\kappa_{1-\alpha,K}:=(1-\alpha)\text{-Quantile}\left(\sup_{\pi\in{K}}{\widehat{Z}}_{\pi}\right)$
$q_{1-\alpha,K}:=(1-\alpha)\text{-Quantile}\left(\sup_{\pi\in{K}}{Z}_{\pi}\right)$
$\widehat{q}_{1-\alpha,K}:=(1-\alpha)\text{-Quantile}^{*}\left(\sup_{\pi\in{K}}{Z}^{*}_{\pi}\right)$
Lemma 1 (Quantile Comparison).
With probability at least $1-\delta_{n}$, critical values obey the following inequalities:
$\kappa_{1-\alpha-2r_{n},K}\leq q_{1-\alpha-r_{n},K}\leq\widehat{q}_{1-\alpha,K}\leq q_{1-\alpha+r_{n},K}$
Furthermore,
$q_{1-\alpha,K}\leq\mathrm{E}\sup_{\pi\in{K}}Z_{\pi}+\sqrt{2\log{1/\alpha}}\leq\sqrt{2\log|K|}+\sqrt{2\log{1/\alpha}}.$
We also note that a lower bound of the form $\sqrt{\log|K|}$ holds, provided that $Z_{\pi}^{\prime}s$ are not too correlated, using Sudakov’s minoration. The upper bound here follows from standard Gaussian concentration bounds.
Proof of Lemma 1.
By Condition (G) we have
$\sup_{{K}\subset\Pi}\sup_{x\in\mathbb{R}}\left|\Pr\left(\sup_{\pi\in{K}}\widehat{Z}_{\pi}\leq x\right)-\Pr\left(\sup_{\pi\in{K}}{Z}_{\pi}\leq x\right)\right|\leq r_{n},$
By Condition (B) we have with probability at least $1-\delta_{n}$:
$\sup_{{K}\subset\Pi}\sup_{x\in\mathbb{R}}\left|{\Pr}^{*}\left(\sup_{\pi\in{K}}\widehat{Z}^{*}_{\pi}\leq x\right)-\Pr\left(\sup_{\pi\in{K}}{Z}_{\pi}\leq x\right)\right|\leq r_{n},$
These relations imply that w.p. $\geq 1-\delta_{n}$:
$(1-\alpha-2r_{n})-\text{Quantile}\left(\sup_{\pi\in{K}}\widehat{Z}_{\pi}\right)\leq(1-\alpha-r_{n})-\text{Quantile}\left(\sup_{\pi\in{K}}{Z}_{\pi}\right)$
$\leq(1-\alpha)\text{-Quantile}^{*}\left(\sup_{\pi\in{K}}{Z}^{*}_{\pi}\right)\leq(1-\alpha+r_{n})\text{-Quantile}\left(\sup_{\pi\in{K}}{Z}_{\pi}\right).$
Thus, the first claim follows.
The second claim follows by the Gaussian concentration inequality of Borell-Sudakov-Tsirelson:
$q_{1-\alpha,K}\leq\mathrm{E}\sup_{\pi\in{K}}Z_{\pi}+\sqrt{2\log{1/\alpha}},$
and by standard calculation we have $\mathrm{E}\sup\limits_{\pi\in{K}}Z_{\pi}\leq\sqrt{2\log|K|}$, where $|{K}|\leq p$ is the cardinality of K{K}. ∎
Proof of Proposition 1
Let
$\Delta^{+}_{\Pi}:=\max_{\pi\in\Pi}(\widehat{Z}_{\pi});\quad\Delta^{-}_{\Pi}:=\min_{\pi\in\Pi}(\widehat{Z}_{\pi});$
[…]
$\underline{\sigma}_{\Pi_{0}}:=\min_{\pi\in\Pi_{0}}\sqrt{n}\widehat{s}(\pi);\quad\widehat{\pi}_{0}=\arg\min_{\pi\in\Pi_{0}}\sqrt{n}\widehat{s}(\pi);$
[…]
We then have
$L(\widehat{k})-V_{\max}$ $=$ $\widehat{V}(\widehat{\pi})-\widehat{k}\widehat{s}(\widehat{\pi})-V_{\max}$
$\geq$ $\widehat{V}(\widehat{\pi}_{0})-\widehat{k}\widehat{s}(\widehat{\pi}_{0})-V_{\max}$
$=$ $V(\widehat{\pi}_{0})+(\widehat{Z}_{\widehat{\pi}_{0}}-\widehat{k})\widehat{s}(\widehat{\pi}_{0})-V_{\max}$
$\geq$ $(\widehat{Z}_{\widehat{\pi}_{0}}-\widehat{k})\widehat{s}(\widehat{\pi}_{0})$
$=$ $(\widehat{Z}_{\widehat{\pi}_{0}}-\widehat{k})\underline{\sigma}_{\Pi_{0}}/\sqrt{n}$
$\geq$ $(\Delta^{-}_{\Pi_{0}}-\widehat{k})\underline{\sigma}_{\Pi_{0}}/\sqrt{n}.$
Also, we have
[…]
$\leq$ $(\widehat{Z}_{\widehat{\pi}}-\widehat{k})_{+}\bar{\sigma}_{\Pi}/\sqrt{n}$
$\leq$ $(\Delta^{+}_{\Pi}-\widehat{k})_{+}\bar{\sigma}_{\Pi}/\sqrt{n}$
Combining the two bounds, we have
[…]
$\leq$ $(-\Delta^{-}_{\Pi_{0}}+\widehat{k})\underline{\sigma}_{\Pi_{0}}/\sqrt{n}+(\Delta^{+}_{\Pi}-\widehat{k})_{+}\bar{\sigma}_{\Pi}/\sqrt{n}$
By the proof of Lemma 1, we have that
$\Pr(\Delta^{+}_{\Pi}\leq q_{1-\beta,\Pi})\geq 1-\beta-r_{n},\quad\Pr(-\Delta^{-}_{\Pi_{0}}\leq q_{1-\beta,\Pi_{0}})\geq 1-\beta-r_{n}.$
It then follows that with probability at least $1-2\beta-2r_{n}$, we have
$V_{\max}-V(\widehat{\pi})\leq(q_{1-\beta,\Pi_{0}}+\widehat{k})\underline{\sigma}_{\Pi_{0}}/\sqrt{n}+(q_{1-\beta,\Pi}-\widehat{k})_{+}\bar{\sigma}_{\Pi}/\sqrt{n}.$
Proof of Proposition 2
Consider the event
$\mathcal{E}_{\Pi}$ $:=$ $\{V(\pi)\geq\widehat{V}(\pi)-\widehat{k}\widehat{s}(\pi),\ \forall\pi\in\Pi\}$
$=$ $\{-\widehat{Z}_{\pi}\geq-\widehat{k},\ \forall\pi\in\Pi\}$
$=$ $\{\max_{\pi\in\Pi}\widehat{Z}_{\pi}\leq\widehat{k}\}$.
By Lemma 1, with our bootstrap choice $\widehat{k}=\widehat{q}_{1-\alpha,\Pi}$ used for $\widehat{\pi}_{\mathrm{PoLeCe}}$, the event
$\mathcal{K}_{\Pi}:=\bigl\{\kappa_{1-\alpha-2r_{n},\Pi}\leq\widehat{k}\leq q_{1-\alpha+r_{n},\Pi}\bigr\}$
holds with probability at least $1-\delta_{n}$. The event $\mathcal{E}_{\Pi}$ occurs with probability at least that of both $\mathcal{E}_{\Pi}$ and $\mathcal{K}_{\Pi}$ jointly. It follows that the event $\mathcal{E}_{\Pi}$ occurs with probability at least $1-\alpha-r_{n}^{\prime}$ for
$r_{n}^{\prime}=2r_{n}+\delta_{n},$
so the claim follows. ∎
Proof of Proposition 3
The first claim was shown in the proof of the previous proposition.
We next show the second claim about the regret bound. The events $\mathcal{E}_{\Pi}$ and $\mathcal{K}_{\Pi}$ from Proposition 2 jointly hold with probability at least $1-\alpha-r_{n}^{\prime}$. Note that $\mathcal{K}_{\Pi}$ implies that $\widehat{q}_{1-\alpha,\Pi}\leq q_{1-\alpha+r_{n},\Pi}$.
Since $V_{\max}=V(\widehat{\pi}_{0})$, under $\mathcal{E}_{\Pi}$ and $\mathcal{K}_{\Pi}$ we have
$V_{\max}-V(\widehat{\pi}_{\mathrm{PoLeCe}})$ $\leq$ $V_{\max}-\max_{\pi\in\Pi}\{\widehat{V}(\pi)-\widehat{q}_{1-\alpha,\Pi}\widehat{s}(\pi)\}$
$\leq$ $V(\widehat{\pi}_{0})-(\widehat{V}(\widehat{\pi}_{0})-\widehat{q}_{1-\alpha,\Pi}\widehat{s}(\widehat{\pi}_{0}))$
$=$ $-\widehat{Z}_{\widehat{\pi}_{0}}\widehat{s}(\widehat{\pi}_{0})+\widehat{q}_{1-\alpha,\Pi}\widehat{s}(\widehat{\pi}_{0})$
$=$ $(\underline{\sigma}_{\Pi_{0}}/\sqrt{n})(\widehat{q}_{1-\alpha,\Pi}-\widehat{Z}_{\widehat{\pi}_{0}})$
$\leq$ $(\underline{\sigma}_{\Pi_{0}}/\sqrt{n})(q_{1-\alpha+r_{n},\Pi}-\Delta^{-}_{\Pi_{0}}).$
By the proof of Lemma 1 we have that $\Pr(-\Delta^{-}_{\Pi}\leq q_{1-\beta,\Pi_{0}})\geq 1-\beta-r_{n}$. Thus
$V_{\max}-V(\widehat{\pi}_{\mathrm{PoLeCe}})\leq V_{\max}-LV_{1-\alpha,\Pi}\leq(\underline{\sigma}_{\Pi_{0}}/\sqrt{n})(q_{1-\alpha+r_{n},\Pi}+q_{1-\beta,\Pi_{0}})$
with probability at least $1-\alpha-\beta-r_{n}-r_{n}^{\prime}$, since we have also used that $\mathcal{E}_{\Pi}$ and $\mathcal{K}_{\Pi}$ both hold, and since
$LV_{1-\alpha,\Pi}=\max_{\pi\in\Pi}\{\widehat{V}(\pi)-\widehat{q}_{1-\alpha,\Pi}\widehat{s}(\pi)\}.$
Appendix B Gaussian Bootstrap for Investment Allocation Problem
The Gaussian bootstrap is a resampling technique widely used in statistical inference to approximate the distribution of a statistic. A critical step in this procedure involves computing the statistic
$\bar{Z}^{*}=\max_{\pi\in\Pi}Z^{*}_{\pi},$
where
$Z^{*}_{\pi}=\frac{\pi^{\top}Z^{*}}{\sqrt{\pi^{\top}\widehat{\Omega}\,\pi}},$
and
$Z^{*}=\widehat{\Omega}^{1/2}\,N^{*}(0,I).$
Here, $N^{*}(0,I)$ denotes a standard Gaussian vector in $\mathbb{R}^{d}$, $\widehat{\Omega}$ is a positive-definite covariance estimator, and $\Pi\subset\mathbb{R}^{d}$ is a convex, bounded subset of the probability simplex.
At first glance, the computation of $\bar{Z}^{*}$ appears to require solving a nonconcave fractional optimization problem:
$\max_{\pi\in\Pi}\frac{\pi^{\top}Z^{*}}{\sqrt{\pi^{\top}\widehat{\Omega}\,\pi}},$
which can be computationally prohibitive. However, by reformulating the problem as a root-finding task involving second-order cone programs, we can achieve computational efficiency.
Reformulation as a Root-Finding Problem
Instead of directly tackling the fractional optimization, we consider the following equivalence:
$\bar{Z}^{*}=\mathrm{root}_{t\geq 0}\left\{\max_{\pi\in\Pi}\left(\pi^{\top}Z^{*}-t\,\sqrt{\pi^{\top}\widehat{\Omega}\,\pi}\right)=0\right\}.$
This reformulation allows us to solve for $\bar{Z}^{*}$ by identifying the unique value of $t$ that zeroes the function
$f(t)=\max_{\pi\in\Pi}\left(\pi^{\top}z-t\,\sqrt{\pi^{\top}\widehat{\Omega}\,\pi}\right)$
when $Z^{*}=z$. Lemma 2, provided in the online supplement, formalizes the properties of $f(t)$ necessary for this approach.
Appendix C Exhibits
(a) Programs targeting age >> 25
(b) Programs targeting age $\leq$ 25
Figure 1. Efficient Decision Frontier.
Notes: Squares represent MVPF estimates and their standard errors for 14 programs collected by the Policy Impacts Library. The red curve shows the decision frontier for budget allocation across these programs. The solid square and circle mark allocations selected by EWM and PoLeCe, respectively. For programs targeting age >> 25, the programs included in the PoLeCe allocation, listed in order of their shares, are: OHIE, Single Adults: 1.16 (0.04), Job Training Partnership Act, Adults: 1.38 (0.40), where each value represents the point estimate (standard error). The program selected by EWM is Holistic Wrap-around Services Can Improve Employment Rates: 1.39 (1.11). For programs targeting age $\leq$ 25, the programs included in the PoLeCe allocation are: Head Start: 1.84 (0.47) and Wisconsin Scholars Grant: 1.43 (0.38). The program selected by EWM is Head Start: 1.84 (0.47).
Table 1. Results for Investment Allocations to Adult Programs (Age >> 25)
EWM PoLeCe
Panel A. Outcome: MVPF
[…]
Panel B. Selected Policy
Holistic Wrap-around Services Can Improve Employment Rates: 1.39 (1.11) 1 0.00
Padua Pilot: 0.18 (0.23) 0 0.00
Housing Vouchers in Chicago: 0.65 (0.02) 0 0.00
Oregon Health Insurance Experiment (Provided to Single Adults): 1.16 (0.04) 0 0.97
Income Maintenance Experiment in Seattle and Denver: 0.27 (0.95) 0 0.00
Paycheck Plus: EITC to Adults without Dependents: 1.00 (0.09) 0 0.00
Work Advance: 0.78 (0.28) 0 0.00
Job Training Partnership Act, Adults: 1.38 (0.40) 0 0.03
Note: EWM denotes empirical welfare maximization, and PoLeCe denotes policy learning with confidence. In Panel A, we report the EWM and the PoLeCe lower confidence bound with $\alpha=0.05$. In Panel B, we report the selected allocation for each program. MVPF estimates and associated standard errors (in parentheses) are included alongside each program name.
Table 2. Results for Investment Allocations to Youth Programs (Age $\leq 25$)
EWM PoLeCe
Panel A. Outcome: MVPF
[…]
Panel B. Selected Policy
Head Start Impact Study: 1.84 (0.47) 1 0.52
Wisconsin Scholar Grant to Low-Income College Students: 1.43 (0.38) 0 0.48
JobStart: 0.20 (0.11) 0 0
Year Up: 0.43 (0.03) 0 0
Job Corps: 0.15 (0.22) 0 0
Job Training Partnership Act, Youth: -0.23 (0.81) 0 0
Note: EWM denotes empirical welfare maximization, and PoLeCe denotes policy learning with confidence. In Panel A, we report the EWM and the PoLeCe lower confidence bound with $\alpha=0.05$. In Panel B, we report the selected allocation for each program. MVPF estimates and associated standard errors (in parentheses) are included alongside each program name.
Online Appendix to “Policy Learning with Confidence”
Victor Chernozhukov, Sokbae Lee, Adam M. Rosen and Liyang Sun
Appendix D Details for Gaussian Bootstrap Implementation
The following Lemma provides formalities for implementing the Gaussian bootstrap procedure described in Appendix B.
Lemma 2 (Existence, Uniqueness, and Differentiability of the Root).
Let $\Pi\subset\mathbb{R}^{d}$ be a convex, closed subset of the probability simplex, and let $\widehat{\Omega}$ be a positive-definite matrix. Define
$f(t)=\max_{\pi\in\Pi}\left(\pi^{\top}z-t\,\sqrt{\pi^{\top}\widehat{\Omega}\,\pi}\right)\quad\text{for }t\geq 0.$
Then:
  1. (1)
    For each fixed $t\geq 0$, the optimization problem
    $\max_{\pi\in\Pi}\left(\pi^{\top}z-t\,\sqrt{\pi^{\top}\widehat{\Omega}\,\pi}\right)$
    can be reformulated as a concave second-order conic optimization problem.
  2. (2)
    There exists a unique $t^{*}\geq 0$ such that $f(t^{*})=0$.
  3. (3)
    The derivative of $f$ at $t$ is given by
    $f^{\prime}(t)=-\,\sqrt{\,\pi^{*}(t)^{\top}\widehat{\Omega}\,\pi^{*}(t)\,}$,
    where $\pi^{*}(t)\in\arg\max_{\pi\in\Pi}\left(\pi^{\top}z-t\,\sqrt{\pi^{\top}\widehat{\Omega}\,\pi}\right)$. Consequently, $f(t)$ is strictly decreasing.
Implications for Gaussian Bootstrap
The reformulation of computing $\bar{Z}^{*}$ as a root-finding problem over a family of second-order cone programs significantly enhances computational efficiency. Each evaluation of $f(t)$ involves solving a concave SOC optimization problem, which is well-supported by modern optimization solvers. The uniqueness of the root $t^{*}$ guarantees that iterative root-finding algorithms, such as the Newton-Raphson method or bisection, will converge reliably to the desired solution. Moreover, the explicit expression for the derivative $f^{\prime}(t)$ facilitates the use of derivative-based optimization methods, further streamlining the computational process.
Proof of Lemma 2.
1. Second-Order Conic Reformulation. Fix any $t\geq 0$. Consider
$f(t)=\max_{\pi\in\Pi}\left(\pi^{\top}z-t\,\sqrt{\pi^{\top}\widehat{\Omega}\,\pi}\right).$
Introduce an auxiliary variable $s$ such that
$s=-\,\sqrt{\pi^{\top}\widehat{\Omega}\,\pi}\quad\Leftrightarrow\quad\|\widehat{\Omega}^{1/2}\,\pi\|_{2}\leq-s,\quad s\leq 0.$
The optimization problem becomes
$\max_{\pi,s}$ $\pi^{\top}z+t\,s$
subject to $\|\widehat{\Omega}^{1/2}\,\pi\|_{2}\leq-s,$
$s\leq 0,$
$\pi\in\Pi.$
The constraint $\|\widehat{\Omega}^{1/2}\,\pi\|_{2}\leq-s$ is a standard second-order cone constraint. Therefore, the problem is a concave second-order conic program.
2. Existence and Uniqueness of the Root. Existence. Observe that
$f(0)=\max_{\pi\in\Pi}\pi^{\top}z,$
which is finite since $\Pi$ is bounded and closed. As $t\to\infty$, the term $t\,\sqrt{\pi^{\top}\widehat{\Omega}\,\pi}$ dominates for every $\pi\in\Pi$, leading to $f(t)\to$-$\infty$. Since $f(t)$ is continuous (as the pointwise maximum of continuous functions over a compact set), the Intermediate Value Theorem guarantees the existence of some $t^{*}\geq 0$ such that $f(t^{*})=0$.
Uniqueness. From part 3, we have
$f^{\prime}(t)=-\,\sqrt{\pi^{*}(t)^{\top}\widehat{\Omega}\,\pi^{*}(t)}<0,$
which implies that $f(t)$ is strictly decreasing. A strictly decreasing continuous function can cross zero at most once. Therefore, the root $t^{*}$ satisfying $f(t^{*})=0$ is unique.
3. Derivative of $f(t)$. Define
$F(\pi,t)=\pi^{\top}z-t\,\sqrt{\pi^{\top}\widehat{\Omega}\,\pi}.$
Then
$f(t)=\max_{\pi\in\Pi}F(\pi,t).$
By the envelope theorem, the derivative of $f(t)$ with respect to $t$ is
$f^{\prime}(t)=\frac{\partial}{\partial t}F(\pi^{*}(t),t)=-\,\sqrt{\pi^{*}(t)^{\top}\widehat{\Omega}\,\pi^{*}(t)}.$
Since $\widehat{\Omega}$ is positive-definite and $\pi^{*}(t)\in\Pi$ with $\Pi$ bounded, it follows that $\pi^{*}(t)^{\top}\widehat{\Omega}\,\pi^{*}(t)>0$. Therefore, $f^{\prime}(t)<0$, establishing that $f(t)$ is strictly decreasing. ∎
Appendix E Alternative Motivations for Risk-Aware Decisions
We consider a decision maker (DM) who holds posterior beliefs $B$ about policy welfare $\{V(\pi)\}$ given by
$V(\pi)\;\sim\;\widehat{V}(\pi)\;+\;\widehat{s}(\pi)\,Z_{\pi},$ (7)
where $\{Z_{\pi}\}$ is a centered Gaussian vector with unit variance $\bigl(\mathrm{E}_{B}[Z_{\pi}^{2}]=1\bigr)$. [^14] This is formally justified by the approximate Bayesian framework of Doksum and Lo (1990), in which the DM treats empirical estimates as the input data and uses approximate likelihoods of these estimators to form beliefs about parameters of interest. Note that $\widehat{V}(\pi)$ and $\widehat{s}(\pi)$ are fixed parameters under the DM’s posterior $B$. They may arise from empirical estimates, incorporate prior information, or reflect other model uncertainties.
The following result shows that this framework encompasses all risk-aware, rational policy choices of a DM whose risk preferences follow the von Neumann-Morgenstern expected utility model. Specifically, suppose the DM has a strictly increasing and concave utility function $U$ and aims to maximize
$\max_{\pi\in\Pi}\,\mathrm{E}_{B}\bigl[U\bigl(V(\pi)\bigr)\bigr].$
Proposition 4 (Posterior Expected Utility Maximization).
Consider the DM described above. Assume $\widehat{V}(\pi)$ and $\widehat{s}(\pi)$ are continuous in $\pi$, and $\Pi$ is compact.
  1. (1)
    If $U$ is strictly concave, then the DM’s choice can be implemented by solving
    $\max_{\pi\in\Pi}\,\Bigl\{\widehat{V}(\pi)\;-\;k_{UB}\,\widehat{s}(\pi)\Bigr\}$ (8)
    for some constant $k_{UB}$ that depends on $U$ and $B$.
  2. (2)
    If $U$ is linear (the DM is risk-neutral), then the DM solves the above program with $k_{UB}=0$. Furthermore, any limit point of the solutions to (8) as $k_{UB}\searrow 0$ lies within the solution set for $k_{UB}=0$.
Thus, a risk-averse DM must solve a variant of (8). As an example, for exponential utility $U(v)=1-\exp(-\lambda v)$, one obtains
$\mathrm{E}[U(V(\pi))]\;=\;1\;-\;\exp\Bigl(-\lambda\,\widehat{V}(\pi)\;+\;\tfrac{1}{2}\,\lambda^{2}\,\widehat{s}^{2}(\pi)\Bigr).$
Maximizing this over $\pi\in\Pi$ is equivalent to (8) for some $k_{UB}>0$. In contrast, setting $k_{UB}=0$ recovers the risk-neutral (EWM) decision. Notably, as $k_{UB}$ approaches zero, solutions to the risk-weighted program (8) converge to the lowest-risk points among the EWM solutions (if the latter is set-valued).
We can gain another perspective by considering a DM with Gaussian beliefs as described above in (7), but whose risk preferences are more directly shaped by a combination of regret and estimation risk, as suggested by Liese and Miescke (2008). [^15] The suggestion was made to unify the problems of selection and estimation.
Consider the loss function
$L(V(\pi))=V_{\max}-V(\pi)+k_{R}\bigl|\widehat{V}(\pi)-V(\pi)\bigr|\sqrt{\bm{\pi}/2},$
where boldface $\bm{\pi}$ denotes the area of the unit circle. This loss function has regret component
$V_{\max}-V(\pi)=\max_{\bar{\pi}\in\Pi}V(\bar{\pi})-V(\pi)$
stemming from using the policy rule $\pi$ rather than the unknown welfare-maximizing choice, and the estimation risk component $\bigl|\widehat{V}(\pi)-V(\pi)\bigr|$. Consider a DM who minimizes the expected regret risk-aware loss function:
$\min_{\pi\in\Pi}\mathrm{E}_{B}\bigl[L(V(\pi))\bigr].$
Proposition 5 (Best Posterior Regret-Risk Aware Decision).
Assume $\widehat{V}(\pi)$ and $\widehat{s}(\pi)$ are continuous in $\pi$, and that $\Pi$ is compact. The DM described above finds the optimal policy by solving:
$\max_{\pi\in\Pi}\,\Bigl\{\,\widehat{V}(\pi)\;-\;k_{R}\,\widehat{s}(\pi)\Bigr\}.$
Propositions 4 and 5 explain why DMs may prefer a strictly positive $k$, but the exact value of $k$ is application-specific. One way to set it is by conducting experiments where DMs respond to hypothetical or incentivized lotteries/choices, then calibrating $k$ based on their observed choices. Another approach is to present a range of efficient decision frontiers, have DMs select their preferred allocations, and reverse-engineer $k$ that aligns with those selections. Generally, there is no single universal way to determine the “right” $k$; its choice depends on context and stakeholder preferences. Using a critical value for $k$ as in PoLeCe provides both interpretable regret and reporting guarantees.
Appendix F Proofs of Propositions 4 and 5
Proof of Proposition 4
Claim (1). Let $v=\widehat{V}(\pi)$ and $s=\widehat{s}(\pi)$. Then
$\mathrm{E}_{B}\bigl[U\bigl(V(\pi)\bigr)\bigr]\;=\;\mathrm{E}\bigl[U(v+sZ)\bigr],\quad Z\sim N(0,1).$
Define $f(v,s):=\mathrm{E}\bigl[U(v+sZ)\bigr].$ The function $(v,s)\mapsto f(v,s)$ is increasing in $v$ and decreasing in $s$. Indeed,
$\frac{\partial}{\partial v}f(v,s)\;=\;\mathrm{E}[U^{\prime}(v+sZ)]\;>\;0,\quad\frac{\partial}{\partial s}f(v,s)\;=\;\mathrm{E}\bigl[U^{\prime}(v+sZ)\,Z\bigr]\;<\;0,$
because $U^{\prime}(\cdot)>0$ and $U^{\prime}(v+sZ)$ is comonotonic with $Z$ (given that $U^{\prime\prime}$ exists by Alexandrov’s theorem almost everywhere and $U^{\prime\prime}<0$ by strict concavity). The second inequality follows from Chebyshev’s association inequality, and the first holds because the expectation of a strictly positive random variable is positive.
[…]
This maximization can be written in a Lagrangian form:
[…]
Claim (2). The first assertion of the claim for $k_{UB}=0$ follows from $\mathrm{E}_{B}[V(\pi)]=\hat{V}(\pi)$. The second assertion of the claim follows from Berge’s maximum theorem. ∎
Proof of Proposition 5.
Observing that
$\mathrm{E}_{B}\bigl[V_{\max}-V(\pi)\bigr]=\mathrm{E}_{B}\bigl[V_{\max}\bigr]-\widehat{V}(\pi),$
and
$\mathrm{E}_{B}\bigl|\widehat{V}(\pi)-V(\pi)\bigr|=\widehat{s}(\pi)\sqrt{\frac{2}{\bm{\pi}}},$
the program becomes
$\min_{\pi\in\Pi}\Bigl\{\mathrm{E}_{B}\bigl[V_{\max}\bigr]-\widehat{V}(\pi)+k_{R}\,\widehat{s}(\pi)\Bigr\}.$
We can now omit the term $\mathrm{E}_{B}\bigl[V_{\max}\bigr]$, resulting in the expression stated in the claim, without affecting the solution of the program.
Appendix G Efficient Decision Frontier Computation
In this appendix, we provide pseudocode of the algorithm to compute the efficient decision frontier. The code below builds a monotonically nondecreasing, concave frontier from the input of points.
Algorithm 1 Build a Monotonically Nondecreasing, Concave Frontier
1:$points=(Risk,Value)$ such that $Risk$ and $Value$ have the same length
2:The set of points that form the efficient frontier
3:procedure FrontierEstimation(pointspoints)
4:  Create data structure from input
5:  Sort by ascending RiskRisk, tie-break by descending ValueValue
6:  $frontier\leftarrow$ empty list
7:  for $i\leftarrow 1$ to the number of rows of pointspoints do
8:   $p\leftarrow\text{row $i$ of $points$}$ $\triangleright$ extract row $i$ from pointspoints
9:   if frontierfrontier is not empty and $p.Value<frontier[\text{last}].Value$ then
10:     continue $\triangleright$ skip to next iteration; ensure monotonicity
11:   end if
12:   while there are at least two elements in frontierfrontier do
13:     $p2\leftarrow\text{last frontier point}$
14:     $p1\leftarrow\text{second to last frontier point}$
15:     if $\frac{p.Value-p1.Value}{p.Risk-p1.Risk}>\frac{p2.Value-p1.Value}{p2.Risk-p1.Risk}$ then
16:      remove $p2$ from frontierfrontier $\triangleright$ ensure concavity
17:     else
18:      break
19:     end if
20:   end while
21:   append $p$ to frontierfrontier
22:  end for
23:  return frontierfrontier $\triangleright$ the set of frontier points
24:end procedure
Note that the algorithm loops over sorted points, skipping any new point with a lower Value than the frontier’s last point, thereby enforcing a monotonically nondecreasing frontier. It also ensures the concavity of the frontier by removing a “bulge,” thus guaranteeing that the slopes decrease as we move right. Here, the bulge is checked by whether the following holds:
$\frac{p.Value-p1.Value}{p.Risk-p1.Risk}\;>\;\frac{p2.Value-p1.Value}{p2.Risk-p1.Risk}.$
That is, each new point to the frontier must produce a slope that is less than or equal to the previous segment’s slope.
For the investment allocation problem, it may be infeasible to discretize $\Pi$ and apply the above algorithm to plot the efficiency frontier. Instead, we start with a fine grid for $k$ centered at $\hat{q}_{0.95,\Pi},$ and directly trace out the efficiency decision frontier by solving the corresponding optimization problem as in  (2). Appendix B provides a computationally efficient way to solve such optimization and therefore to plot the efficiency frontier.
Appendix H Treatment Choice
This appendix considers the problem of selecting treatment choice policies. Section H.1 describes the treatment choice problem, and Section H.2 then demonstrates application of PoLeCe in three empirical settings using data from RCTs.
H.1. Problem of Choosing Better Treatment Policies.
Given a finite treatment set $\mathcal{T}$ and potential outcomes $Y(t)$ for each treatment $t$, a policy $\pi\in\Pi$ specifies probabilities $\pi(t\mid X)$ of assigning treatment $t$ given covariates $X$, where $\Pi$ denotes the set of allowable policies. For example, if any fractional allocation is allowed $\Pi$ is the unit simplex, while if treatment assignment is required to be a deterministic function of covariates then $\Pi$ is the set of vertices of the unit simplex. Other constraints can also be accommodated by appropriate specification of $\Pi$.
The welfare function of a policy is then defined as
$V(\pi)\;=\;\mathrm{E}\!\Bigl[\sum_{t\in\mathcal{T}}Y(t)\,\pi\bigl(t\mid X\bigr)\Bigr],$
which captures the expected outcome under policy $\pi$.
We assume we have data $(W_{i})_{i=1}^{n}$ in the form of i.i.d. copies of $W=(Y,T,X)$, where $T$ is the treatment, $X$ are covariates, and $Y$ is the outcome. We also assume the data were collected under the unconfoundedness assumption, namely that the assigned treatment $T$ is independent of the potential outcome $Y(t)$ given $X$. This implies that $\mathrm{E}[Y(t)\mid X]$ is identified from the regression
$g(t,X)=\mathrm{E}[Y\mid T=t,X],$
provided that $p(t\mid X)=\Pr[T=t\mid X]>0$.
Let $p(t\mid X)=\Pr[T=t\mid X]$ be the propensity score, and define the Riesz representer
$H(\pi)\;=\;\sum_{t\in\mathcal{T}}\mathbf{1}\{T=t\}\,\frac{\pi(t\mid X)}{p(t\mid X)}.$
Under standard conditions, including $\mathrm{E}[H^{2}(\pi)]<\infty$ – a relevant overlap condition [^16] This requires that policies do not place weight on treatments for which there is insufficient randomization at particular values of $X$. Note that the standard overlap condition can fail, but the welfare of a particular policy is identified, for example, if $H(\pi)$ and $Y$ have a finite second moment. – we obtain the following regression and representer formulas for identifying the policy’s welfare:
$V(\pi)\;=\;\mathrm{E}\biggl[\sum_{t\in\mathcal{T}}g(t,X)\,\pi(t\mid X)\biggr]\;=\;\mathrm{E}\!\bigl[Y\,H(\pi)\bigr],$ (9)
which can be viewed either as aggregating up regressions or as reweighting observed outcomes by how the policy $\pi$ differs from the propensity score $p$. Combining these two approaches leads to efficient representation and estimation of the welfare function via the influence function (e.g., Newey, 1994):
$V(\pi)\;=\;\mathrm{E}[\psi_{\pi}(W)],\quad\text{where}\quad\psi_{\pi}(W)\;:=\;\sum_{t\in\mathcal{T}}g(t,X)\,\pi(t\mid X)\;+\;H(\pi)\,\bigl[Y-g(T,X)\bigr],$
also known as the doubly-robust score in a parametric estimation context.
Added Value of Policies vs. Control Policy. It is often the case that we want to make decisions depending upon whether policies improve upon a control treatment. Let us denote by $t=0$ the control state. In this case, we can re-define:
$V(\pi)\;=\;\mathrm{E}\!\Bigl[\sum_{t\in\mathcal{T}}\{Y(t)-Y(0)\}\,\pi\bigl(t\mid X\bigr)\Bigr].$ (10)
The efficient score for estimation and inference for $V(\pi)$ is now
$\psi_{\pi}(W):=\sum_{t\in\mathcal{T}}(g(t,X)-g(0,X))\pi(t\mid X)+H(\pi)\,\bigl[Y-g(T,X)\bigr]$ (11)
where we re-define $H(\pi)=\sum_{t\in\mathcal{T}}\ \left[\frac{\mathbf{1}\{T=t\}}{p(t\mid X)}-\frac{\mathbf{1}\{T=0\}}{p(0\mid X)}\right]\pi(t\mid X)$.
We follow the convention of encapsulating in $X$ both covariates that are confounders as well as those for which the planner may assign different treatments.
In randomized experiments, the representers are known, and we only need consistent cross-fitted learners for the regression functions to obtain consistent, asymptotically normal, efficient estimators $\widehat{V}(\pi)$. [^17] With randomized experimental data, the pseudo-consistency of a regression estimator for some function (though not necessarily the true function) is sufficient to satisfy all desirable properties except for asymptotic efficiency. In observational studies, one can employ cross-fitted, modern statistical or machine learning methods to estimate the Riesz representers and the regressions. The resulting estimators satisfy
$\widehat{V}(\pi)=V(\pi)+\mathbb{E}_{n}\bigl(\psi_{\pi}(W_{i})-V(\pi)\bigr)+o_{p}\bigl(1/\sqrt{n}\bigr),$
uniformly over $\pi\in\Pi$, where $\Pi$ is either a finite set or a set whose complexity does not grow too fast. By further applying high-dimensional central limit theorems, we obtain
$\{\widehat{V}(\pi)\}\;\overset{a}{\sim}\;\{V(\pi)+s(\pi)\,Z_{\pi}\},\quad Z_{\pi}\sim N(0,1),\quad s^{2}(\pi)=\mathrm{Var}(\psi_{\pi})/n,$
where $\{Z_{\pi}\}_{\pi\in\Pi}$ is a centered Gaussian process, with each component having standard normal distribution; and $\overset{a}{\sim}$ means “approximately distributed as” in the sense of Condition (G) in Section 2.2.
H.2. Empirical Applications to Treatment Choice
In this section we demonstrate application of PoLeCe to the problem of choosing better treatment policies. We use RCT data from Banerjee et al. (2025), Dupas and Robinson (2013), and Schilbach (2019) to illustrate the usefulness of our method and to show how the results can be incorporated into research papers that analyze RCTs. These papers focus on estimating the added value relative to control (as defined in (10)). In contrast, we focus on each policy’s value to fully account for the estimation error for estimating the average outcome under control. While these papers originally used regression adjustment, we rely on efficient influence functions (doubly robust scores via (11)) to obtain welfare estimates for potential precision gains.
H.2.1. Immunization Nudges
The first example illustrates the trade-off between welfare and precision in selecting the best arm. Banerjee et al. (2025) conducted a large-scale RCT of nudges to encourage immunization in the state of Haryana, Northern India. The study employed a cross-randomized design involving three main types of nudges: providing incentives, sending short messaging service (SMS) reminders, and seeding community ambassadors. Each type included multiple variations, resulting in 74 distinct treatments and a control. While Banerjee et al. (2025) proposed a method to group similar treatments for analysis, we retain the original set of treatments for illustrative purposes.
The outcomes include the number of measles shots administered, and number of shots per dollar spent. We observe 7,370 households and use doubly robust scores to obtain the welfare estimates $\widehat{V}(\pi)$ and their standard errors. Following Banerjee et al. (2025), we weight these village-level regressions by village population, and standard errors are clustered at the SC level. We calculate $\widehat{k}=\widehat{q}_{1-\alpha,\Pi}$ by the bootstrap procedure in Condition (B).
(a) Measles shots outcome
(b) Shots per dollar outcome
Figure 2. Welfare estimates against precision
Notes: Each blue circle represents a pair of welfare estimate and its standard error $(\widehat{V}(\pi),\widehat{s}(\pi))$ for a given treatment $\pi$. The connected red circles show the decision frontier.
As shown in Figure 2 , there is substantial variation in the precision of welfare estimates across treatments. While EWM selects the treatment with the highest welfare estimate $\widehat{V}(\pi)$, PoLeCe adjusts for estimation uncertainty and may choose differently. Here we set $\alpha=0.05$ which places large emphasis on precision. For example, for shots per dollar, EWM and PoLeCe both select trusted info hub, no incentive, no reminder, as there is minimal trade-off between welfare and precision. However, for number of shots, EWM selects trusted info hub, high slope, high SMS, while PoLeCe selects info hub, low slope, low SMS because the latter has comparable welfare but much higher precision.
Table 3. Results for Immunization Nudges
EWM PoLeCe Control
Panel A. Outcome: Measles shots
[…]
Panel B. Outcome: Shots per dollar
[…]
Panel C. Selected Policy:
Measles shots trusted info hub, info hub, no info,
high slope low slope no incentive
high SMS low SMS no reminder
Shots per dollar trusted info hub, trusted info hub, no info,
no incentive no incentive no incentive
no reminder no reminder no reminder
Note. EWM refers to empirical welfare maximization, and PoLeCe corresponds to policy learning with confidence. In Panels A and B, the EWM and PoLeCe lower confidence bounds with $\alpha=0.05$ are provided. In Panel C, selected treatment combinations are given.
H.2.2. Informal Savings Technologies
Dupas and Robinson (2013) conducted a field experiment in Kenya to study how simple informal savings technologies can increase investment in preventative health and reduce vulnerability to health shocks. The experiment was run through Rotating Savings and Credit Associations (ROSCAs), and all participants were required to be enrolled in a ROSCA at the start. [^18] ROSCAs are informal savings groups. Members come together on a regular basis and contribute to a common pot of money which is taken home by one member on a rotating basis.
Specifically, 771 ROSCA participants were randomized to five treatment groups: (i) Control; (ii) Safe Box: participants were given a simple locked box made out of metal, while they were asked to record what health product they were saving for, and its cost, on a passbook; (iii) Lockbox: participants were given a passbook and a locked box identical to those in the Safe Box treatment, except that the program officer kept the key; (iv) Health Pot: participants were given a side pot that the members could contribute to in addition to the regular ROSCA pot; [^19] Unlike the regular pot, this pot would be earmarked for a specific health product. (v) Health Savings Account, or HSA: participants were encouraged to make regular deposits into an individual HSA managed by the ROSCA treasurer.
The outcomes include investments in health and measures of whether people have trouble affording medical treatments. As an illustration, we focus on health investments in terms of amount (in Ksh) spent on preventative health products. We choose two baseline covariates: Female and Married purely for simplicity of illustration, without taking a stand on whether they should be used in this setting. [^20] In practice, equity, fairness or other concerns could make the use of these covariates for heterogeneous policy assignment unwarranted, and the use of other covariates, either in addition or separately, may be preferred. One could then compute the PoLeCe rule that incorporates the desired covariates. We obtain 691 individuals after removing 80 participants who received multiple treatments. The randomization was done after stratifying on some ROSCA characteristics. We incorporate strata indicators as well as Female and Married for calculating the doubly robust scores.
Table 4. Results for Informal Savings Technologies
EWM PoLeCe Control
Panel A. Outcome: Health Investment (in Ksh)
[…]
Panel B. Selected Policy:
(Female, Married) =(0,0)=(0,0) Health Pot Health Pot Control
(Female, Married) =(0,1)=(0,1) Health Pot Health Pot Control
(Female, Married) =(1,0)=(1,0) Health Pot Safe Box Control
(Female, Married) =(1,1)=(1,1) Safe Box Health Pot Control
Note. EWM refers to empirical welfare maximization, and PoLeCe corresponds to policy learning with confidence. In Panel A, EWM and PoLeCe lower confidence bounds with $\alpha=0.05$ are provided along with that for the control group policy. In Panel B, selected treatment for each vector of (Female, Married) is given.
There are $|\Pi|=625=5^{4}$ possible treatment policies all together because there are five treatment values and four possible values in the support of the two covariates. Figure 3(a) plots all welfare estimates relative to their precision. The highest welfare is observed below 900 Ksh, but it is accompanied by the largest standard error. Therefore, it is advisable to account for the precision of these estimates when selecting an optimal policy. To implement PoLeCe at the level $\alpha=0.05$ and calculate $\widehat{q}_{0.95,\Pi}$, we apply the bootstrap procedure from Condition (B). Table 4 summarizes the empirical results. In Panel A, the maximized welfare by EWM is 890 Ksh, which corresponds to the highest welfare in Figure 3(a). The adjusted estimate of maximum welfare $\widehat{V}(\widehat{\pi}_{\mathrm{PoLeCe}})$ at $\alpha=0.05$ is 587, which is much less than no-precision-corrected $\widehat{V}(\widehat{\pi}_{\text{EWM}})$ but substantially greater than $\widehat{V}(\widehat{\pi}_{\text{EWM}})-\widehat{q}_{0.95,\Pi}\widehat{s}(\widehat{\pi}_{\text{EWM}})$. Furthermore, the chosen optimal policies differ for females. EWM selects Health Pot for unmarried women and Safe Box for married women. In contrast, after accounting for standard errors, PoLeCe recommends Safe Box for unmarried women and Health Pot for married women.
(a) Health investment (in Ksh)
(b) Labor supply
Figure 3. Welfare estimates against precision
Notes: Each blue circle represents a pair of welfare estimate and its standard error $(\widehat{V}(\pi),\widehat{s}(\pi))$ for a given treatment $\pi$. The connected red circles show the decision frontier.
H.2.3. Alcohol and Self-Control
Schilbach (2019) studied alcohol consumption among low-income workers in India. Specifically, 229 cycle-rickshaw drivers in India were randomized to three treatment groups: (i) control group: participants were paid Rs 90 (\$1.50) each day for visiting the study office; (ii) incentive group: drivers were given incentives to remain sober—the payment was Rs 60 (\$1.00) if they arrived at the office with a positive blood alcohol content (BAC) and Rs 120 (\$2.00) if they arrived sober; (iii) choice group: individuals were given the same incentives earlier in the study and then offered to choose between incentives and unconditional payments in the later phase of the study. See Schilbach (2019) for details on the experiment and background.
As an illustration, we focus on labor supply measured by the fraction of days individuals worked from day 5 (the first day of sobriety incentives) through day 19 (the last day of sobriety incentives). For simplicity, we choose two baseline covariates: Baseline sober =1=1 if the baseline fraction sober is 1 and Baseline sober =0=0 otherwise; Owns rickshaw =1=1 if a driver own a rickshaw and Owns rickshaw =0=0 otherwise. After removing observations with missing values, there are 222 individuals. Given complete randomization, we calculate doubly robust scores based on the two baseline covariates.
Table 5. Results for Self-control
EWM PoLeCe Control
Panel A. Outcome: Labor Supply
[…]
Panel B. Selected Policy:
(Baseline sober, Owns rickshaw) =(0,0)=(0,0) Control Control Control
(Baseline sober, Owns rickshaw) =(0,1)=(0,1) Control Incentive Control
(Baseline sober, Owns rickshaw) =(1,0)=(1,0) Choice Choice Control
(Baseline sober, Owns rickshaw) =(1,1)=(1,1) Choice Choice Control
Note. EWM refers to empirical welfare maximization, and PoLeCe corresponds to policy learning with confidence. In Panel A, the EWM point estimate and the PoLeCe with $\alpha=0.05$ are provided along the control group policy. In Panel B, selected treatment for each vector of (Baseline sober, Owns rickshaw) is given.
There are $|\Pi|=81=3^{4}$ possible treatment policies in total since there are three treatment levels and four possible values of the two covariates. Figure 3(b) plots the welfare estimates against their precision. Although the highest welfare appears with a relatively small standard error, this does not necessarily imply statistical significance. To implement PoLeCe at the level $\alpha=0.05$ and calculate $\widehat{q}_{0.95,\Pi}$, we apply the bootstrap procedure from Condition (B). Table 5 summarizes the empirical results. The EWM and PoLeCe estimates, along with their lower confidence bounds, are very similar and only slightly exceed those of the control group.
Appendix I Computational Experiments
Here we provide results of computational experiments calibrated to the three empirical applications in Section H. Specifically, we take the welfare estimates and their standard errors $(\widehat{V}(\pi),\widehat{s}(\pi))$ as shown in Figure 2 (DGP 1 and 2), Figure 3(a) (DGP 3) and Figure 3(b) (DGP 4) as true parameter values $(V(\pi),s(\pi))$. For each DGP, we take independent draws from $N(V(\pi),s^{2}(\pi))$ for $\pi\in\Pi$ and then assess the performance of EWM and PoLeCe based on $(\widehat{V}^{\ast}(\pi),s(\pi))$ where $\widehat{V}^{\ast}(\pi)$ are the simulation draws. We focus on $\alpha=0.05$ and since we take $\widehat{V}^{\ast}(\pi)$ to be independent across $\pi\in\Pi$, we obtain the critical value $\widehat{q}_{1-\alpha,\Pi}$ via the bootstrap as discussed in Condition (B) by setting the correlation matrix to be the identity matrix. As a benchmark, we also report the performance of choosing the control policy, whose identity is known in each experiment.
Table 6 summarizes the simulation results. In line with our theoretical results, when there is heterogeneity in the estimation risk for policies and $\underline{\sigma}_{\Pi_{0}}$ (the lower bound on the risk of the best policies $\Pi_{0}$) is small, as in DGP 1, 2 and 4, PoLeCe outperforms EWM in terms of tail regret. Since PoLeCe is data-dependent, there is still variability in regret, but it only loses to the control policy in rare cases, as shown in DGP 4. Finally, by construction, the lower confidence band for PoLeCe is the highest, and is much higher than EWM in all four DGPs considered in this simulation.
Table 6. Calibrated Simulation Results
Average Median 95%-percentile Average
regret regret regret Welfare LCB
Measles shots EWM 7.88% 2.12% 32.10% -36.20%
(DGP 1) PoLeCe 8.83% 10.02% 12.84% -21.53%
Control 52.64% 52.64% 52.64% -79.92%
Shots per dollar EWM 1.74% 1.77% 4.89% -6.64%
(DGP 2) PoLeCe 1.00% 0.00% 3.83% -5.98%
Control 4.89% 4.89% 4.89% -16.31%
Informal Saving Technology EWM 15.91% 17.21% 33.85% -54.17%
(DGP 3) PoLeCe 28.14% 30.77% 41.45% -45.02%
Control 66.39% 66.39% 66.39% -87.82%
Alcohol and Self-Control EWM 1.38% 1.08% 4.46% -3.66%
(DGP 4) PoLeCe 0.84% 0.47% 2.21% -3.46%
Control 1.43% 1.43% 1.43% -7.86%
[…]